# Leveraging Large Language Models for Code Translation and Software Development in Scientific Computing

### v1

@@ Abstract.
The emergence of foundational models and generative artificial intelligence (GenAI) is poised to transform productivity in scientific computing, especially in code development, refactoring, and translating from one programming language to another. However, because the output of GenAI cannot be guaranteed to be correct, manual intervention remains necessary. Some of this intervention can be automated through task-specific tools, alongside additional methodologies for correctness verification and effective prompt development. We explored the application of GenAI in assisting with code translation, language interoperability, and codebase inspection within a legacy Fortran codebase used to simulate particle interactions at the Large Hadron Collider (LHC). In the process, we developed a tool, [Code-Scribe] , which combines prompt engineering with user supervision to establish an efficient process for code conversion. In this paper, we demonstrate how [Code-Scribe] assists in converting Fortran code to C++, generating Fortran-C APIs for integrating legacy systems with modern C++ libraries, and providing developer support for code organization and algorithm implementation. We also address the challenges of AI-driven code translation and highlight its benefits for enhancing productivity in scientific computing workflows.
@@ 1. Introduction
In practice, manual incremental translation has been the predominant approach in the community. Typically, developers convert code sections gradually, build Fortran-C interfaces, and run tests to ensure correctness. This controlled approach allows developers to verify the functionality of each section before moving on to the next. However, this process is labor-intensive, requiring extensive knowledge of both Fortran and C++ and an understanding of the interoperability challenges between the two languages. We investigated the use of GenAI to aid in translating MCFM , a Monte Carlo code that simulates particle interactions observed at the Large Hadron Collider (LHC), from Fortran to C++. To this end, we developed a tool, [Code-Scribe] , along with a methodology for efficient translation. This paper describes [Code-Scribe] and its application to MCFM and also highlights other legacy Fortran codes under consideration.
The paper is organized as follows: Section 2 reviews the literature on code translation and the use of GenAI for developer productivity. Section 3 provides an overview of [Code-Scribe] , followed by the translation methodology in Section 4. Section 5 presents results from large language model sensitivity studies, and Section 6 discusses ongoing and future work.
@@ 2. Background
We have developed [Code-Scribe] (Dhruv, 2024a), an AI-assisted tool designed to facilitate the translation of Fortran codebases to C++ as a step towards integrating GenAI technology in scientific code development workflows. We explore its application in code translation for scientific computing, and compare the performance and correctness of different LLMs in this context. Additionally, we highlight how [Code-Scribe] has been applied to various scientific computing software projects, for code translation and code inspection to enhance developer productivity. Finally, we outline future directions for enhancing the tool and integrating it into scientific workflows.
3. Overview of [Code-Scribe]
In scientific computing, the urgency to translate legacy Fortran codebases to modern languages like C++ is driven by the desire to leverage advanced libraries and hardware optimizations for HPC platforms. While Fortran remains highly performant, its support for platform heterogeneity remains subpar compared to that of C++. Translating legacy codes to C++ makes a much richer ecosystem for performance portability accessible to these codes. However, this translation process is fraught with challenges.
@@ 3. Overview of [Code-Scribe]
Despite its advantages, incremental translation remains labor-intensive. Developers must manually manage Fortran-C layers, handle complex data structures, and ensure that performance optimizations are retained. This is where GenAI can make a significant impact. However, effectively leveraging AI within an application requires a careful understanding of the task at hand, domain insight, and effective prompt development. Through our exploration of this field, we distilled our knowledge into [Code-Scribe] , a tool that offers both code translation and software development assistance.
@@
Schematic of the workflow required for a systematic code conversion process. Steps in blue are managed using [Code-Scribe] while steps in red are manual and require developer intervention
[Code-Scribe] operates within a systematic framework that uses LLMs to aid developers in managing the complexities of code conversion. It uses chat completion as a prompting technique to generate desired coding output from GenAI models. This interaction-based strategy achieves good results through structured conversations that guide the code generation process. This interaction forms the foundation of [Code-Scribe] ’s methodology, creating templates to streamline AI-assisted code conversion.
@@
Initially, the tool maps the project structure by indexing subroutines, modules, and functions across various files, which is critical for understanding the relationships between components in the codebase. Once the project structure is mapped, [Code-Scribe] generates a draft of the C++ code for a given Fortran source file. It employs pattern recognition to replace variable declarations, include appropriate header files, and identify the use of external modules and subroutines. This information is encoded in the draft to help LLMs better understand the code context. Using the chat template as a seed prompt, [Code-Scribe] appends the source and draft files and triggers the LLM to complete the code conversion. The generated results are extracted, reviewed, compiled, and tested for correctness. Any errors found during compilation are addressed manually by the developer or sent for regeneration by updating the seed prompt. While [Code-Scribe] automates many aspects of the translation process, human expertise remains essential for final review and correctness, ensuring the translated code adheres to scientific computing standards.
Figure 3 provides an overview of the workflow implemented by [Code-Scribe] . The schematic illustrates the systematic code conversion process, distinguishing between automated and manual steps. Blue steps indicate tasks handled by [Code-Scribe] , such as project indexing, draft generation, and LLM-driven translations, while red steps indicate manual activities, including chat template generation, code review, and verification.
@@ 4. Methodology
Based on the overview in the previous section and the schematic in Figure 3, we developed a command-line interface for [Code-Scribe] with four commands: (1) Index, (2) Inspect, (3) Draft, and (4) Translate. Each command incorporates elements of RAG, enhancing the effectiveness of the translation process. Detailed documentation for the usage of these commands can be found in the [Code-Scribe] repository (Dhruv, 2024a); here, we focus on providing an overview of their functionality.
@@ 4.1. Index
This command generates a comprehensive mapping of the source tree by analyzing the project’s file hierarchy, dependencies, and code relationships. It creates a YAML file within each subdirectory, capturing critical metadata such as the relative directory location, file names, and associated constructs like modules, subroutines, and functions. This structured index is essential for navigating large legacy codebases and ensures that subsequent translation steps are both efficient and accurate. The design of the Index command embodies RAG principles by enabling efficient retrieval of relevant metadata, which aids in informed querying during later commands. Figure 4 illustrates an example of the YAML index file. The file functions as a dictionary, where the key directory stores the relative path of a subdirectory to the project root defined by the key root. The key files contains a hierarchy of dictionaries with keys corresponding to filenames and constructs. For example: [files][file1.f90][modules] provides a list of modules in the file file1.f90. During the execution of other [Code-Scribe] commands, YAML files from all directories are loaded and compiled into an inverse dictionary that maps constructs back to their corresponding filenames for efficient querying. For instance, during draft code generation, the inverse dictionary is utilized to determine that subroutineA belongs to file2.f90.
@@ 4.4. Translate
The core functionality of [Code-Scribe] lies in its AI-powered translation. This command transforms Fortran source code into C++ while generating the necessary Fortran-C interfaces for smooth integration between legacy and modern codebases. The translation leverages generative AI to interpret Fortran’s syntax and semantics. The seed prompt from the chat template is appended with code from the source and draft files and sent to the AI for chat completion. The source code from the Fortran file is supplied within the elements <source>...</source>, and the draft code is enclosed within <draft>...</draft>. The LLM is instructed to produce output for the C++ source and the corresponding Fortran-C interface, enclosing them within <csource>...</csource> and <fsource>...</fsource>. This mechanism is based on the implementation by the Slack team in their recent article (Slack, 2024), allowing for source code extraction and saving to separate files named <filename>.cpp and <filename>_fi.f90.
@@
5. LLM Sensitivity Studies
We explored the application of [Code-Scribe] to the following scenarios within scientific computing:
@@ 5. LLM Sensitivity Studies
Developing C++ interface for Noah-MP: Noah-MP (He, C., et al., 2023) is a Fortran library that models urban canopy parameterization for atmospheric simulations. Integrating it with the Energy Research and Forecasting (ERF) Model (Almgren et al., 2023) is essential for enhancing the latter’s applicability. However, ERF is written in C++. We utilized [Code-Scribe] to create interfaces for this integration.
The discussion is dominated by MCFM code conversion because that was not only the motivator for the development of [Code-Scribe] but has seen its most extensive use. As of this writing the entire code has been converted to C++. The conversion process initially averaged 2-3 files per day before the development of the tool. After the tool was implemented, developer productivity increased to 10-12 files per day. Below are a few additional statistics gleaned from this work.
@@ 5.1. Model Sensitivity Study
For this study we focused on evaluating the performance of various GenAI models and provide a quantitative assessment of the developer time required to verify and test the code after LLM conversion. The raw source files used in this analysis are available in (Dhruv, 2024b), and for the rest of this discussion, we will refer to the data repository as <results>.
@@
GPT-4o and GPT-3.5 Turbo, developed by OpenAI, are general-purpose models accessible via API endpoints, requiring the purchase of API credits. GPT-4o is a multimodal model capable of processing both text and images, and although OpenAI has not disclosed its architecture or parameter count, it is widely speculated to have significantly more parameters than other models. GPT-3.5 Turbo is smaller and optimized for cost-effective, high-speed text generation. While GPT-4o is well-suited for complex multimodal tasks and excels in multi-step reasoning, CodeLlama and Mistral are specialized for coding tasks and do not support multi-step reasoning to the same extent as GPT-4o. This distinction is critical when evaluating model performance in specific use cases, such as the MCFM code conversion, where the ability to engage in multi-step reasoning can significantly enhance the translation process.
@@ 5.2. RAG Sensitivity Study
In this study, we focus on how retrieval-augmented generation (RAG) enhances code translation performance. RAG has previously been shown to improve the accuracy of natural language processing (NLP) tasks, as demonstrated by Lewis et al. (Lewis et al., 2020). To evaluate its effectiveness in code translation, we applied RAG elements within [Code-Scribe] , specifically during the Draft command. In this step, a target.scribe file is created, and RAG is employed by parsing YAML index files and annotating the draft with prompts that provide context to the LLM, particularly about the presence of external functions.
@@ 6. Ongoing Work
In our ongoing work, we are applying the code conversion techniques developed for MCFM to design a Fortran-C interoperability layer for Noah-MP and ERF, with the goal of incorporating Urban Canopy parameterization into ERF simulations. Additionally, we are leveraging a RAG-based approach to develop modules for ERF, utilizing information from the existing codebase to seamlessly integrate new physics with the overall system. This approach will introduce a new functionality called Generate into [Code-Scribe] , which will assist in writing new modules by drawing on context from the surrounding code.
To enhance developer productivity during code conversion, we plan to expand the Inspect command to automatically generate seed_prompt.toml chat templates and to integrate [Code-Scribe] with the LASSI framework (Dearing et al., 2024) to improve the correctness of the generated code.
@@ 7. Conclusions
In this study we presented [Code-Scribe] , a GenAI-driven tool designed to facilitate the translation of legacy Fortran code to modern C++ while creating necessary interfaces for seamless integration. Our evaluations demonstrate that [Code-Scribe] reduces the developer time required for code conversion and testing, showcasing its potential to enhance productivity in scientific computing projects.
@@
In addition to the MCFM code conversion, our ongoing efforts focus on applying [Code-Scribe] to meet developer requirements across various scientific computing projects, such as Noah-MP and Flash-X, by enhancing existing commands and introducing new functionality.
As we continue to refine [Code-Scribe] and explore its applications, our aim is to further streamline the translation process and reduce overhead associated with code review and testing by exploring its integration with the LASSI framework developed at Argonne. We envision [Code-Scribe] as a valuable tool that empowers developers in scientific computing to leverage GenAI effectively.

### v2

@@ Abstract.
The emergence of foundational models and generative artificial intelligence (GenAI) is poised to transform productivity in scientific computing, especially in code development, refactoring, and translating from one programming language to another. However, because the output of GenAI cannot be guaranteed to be correct, manual intervention remains necessary. Some of this intervention can be automated through task-specific tools, alongside additional methodologies for correctness verification and effective prompt development. We explored the application of GenAI in assisting with code translation, language interoperability, and codebase inspection within a legacy Fortran codebase used to simulate particle interactions at the Large Hadron Collider (LHC). In the process, we developed a tool, CodeScribe, which combines prompt engineering with user supervision to establish an efficient process for code conversion. In this paper, we demonstrate how CodeScribe assists in converting Fortran code to C++, generating Fortran-C APIs for integrating legacy systems with modern C++ libraries, and providing developer support for code organization and algorithm implementation. We also address the challenges of AI-driven code translation and highlight its benefits for enhancing productivity in scientific computing workflows.
@@ 1. Introduction
In practice, manual incremental translation has been the predominant approach in the community. Typically, developers convert code sections gradually, build Fortran-C interfaces, and run tests to ensure correctness. This controlled approach allows developers to verify the functionality of each section before moving on to the next. However, this process is labor-intensive, requiring extensive knowledge of both Fortran and C++ and an understanding of the interoperability challenges between the two languages. We investigated the use of GenAI to aid in translating MCFM, a Monte Carlo code that simulates particle interactions observed at the Large Hadron Collider (LHC), from Fortran to C++. To this end, we developed a tool, CodeScribe, along with a methodology for efficient translation. This paper describes CodeScribe and its application to MCFM and also highlights other legacy Fortran codes under consideration.
The paper is organized as follows: Section 2 reviews the literature on code translation and the use of GenAI for developer productivity. Section 3 provides an overview of CodeScribe, followed by the translation methodology in Section 4. Section 6 presents results from large language model sensitivity studies, and Section 7 discusses ongoing and future work.
@@ 2. Background
We have developed CodeScribe (Dhruv, 2024a), an AI-assisted tool designed to facilitate the translation of Fortran codebases to C++ as a step towards integrating GenAI technology in scientific code development workflows. We explore its application in code translation for scientific computing, and compare the performance and correctness of different LLMs in this context. Additionally, we highlight how CodeScribe has been applied to various scientific computing software projects, for code translation and code inspection to enhance developer productivity. Finally, we outline future directions for enhancing the tool and integrating it into scientific workflows.
3. Overview of CodeScribe
In scientific computing, the urgency to translate legacy Fortran codebases to modern languages like C++ is driven by the desire to leverage advanced libraries and hardware optimizations for HPC platforms. While Fortran remains highly performant, its support for platform heterogeneity remains subpar compared to that of C++. Translating legacy codes to C++ makes a much richer ecosystem for performance portability accessible to these codes. However, this translation process is fraught with challenges (Cary et al., 1997).
@@ 3. Overview of CodeScribe
Despite its advantages, incremental translation remains labor-intensive. Developers must manually manage Fortran-C layers, handle complex data structures, and ensure that performance optimizations are retained. This is where GenAI can make a significant impact. However, effectively leveraging AI within an application requires a careful understanding of the task at hand, domain insight, and effective prompt development. Through our exploration of this field, we distilled our knowledge into CodeScribe, a tool that offers both code translation and software development assistance.
@@
Schematic of the workflow required for a systematic code conversion process. Steps in blue are managed using CodeScribe while steps in red are manual and require developer intervention
CodeScribe operates within a systematic framework that uses LLMs to aid developers in managing the complexities of code conversion. It uses chat completion as a prompting technique to generate desired coding output from GenAI models. This interaction-based strategy achieves good results through structured conversations that guide the code generation process. This interaction forms the foundation of CodeScribe’s methodology, creating templates to streamline AI-assisted code conversion.
@@
Initially, the tool maps the project structure by indexing subroutines, modules, and functions across various files, which is critical for understanding the relationships between components in the codebase. Once the project structure is mapped, CodeScribe generates a draft of the C++ code for a given Fortran source file. It employs pattern recognition to replace variable declarations, include appropriate header files, and identify the use of external modules and subroutines. This information is encoded in the draft to help LLMs better understand the code context. Using the chat template as a seed prompt, CodeScribe appends the source and draft files and triggers the LLM to complete the code conversion. The generated results are extracted, reviewed, compiled, and tested for correctness. Any errors found during compilation are addressed manually by the developer or sent for regeneration by updating the seed prompt. While CodeScribe automates many aspects of the translation process, human expertise remains essential for final review and correctness, ensuring the translated code adheres to scientific computing standards.
Figure 3 provides an overview of the workflow implemented by CodeScribe. The schematic illustrates the systematic code conversion process, distinguishing between automated and manual steps. Blue steps indicate tasks handled by CodeScribe, such as project indexing, draft generation, and LLM-driven translations, while red steps indicate manual activities, including chat template generation, code review, and verification.
@@ 4. Methodology
Based on the overview in the previous section and the schematic in Figure 3, we developed a command-line interface for CodeScribe with four commands: (1) Index, (2) Inspect, (3) Draft, and (4) Translate. Each command incorporates elements of RAG, enhancing the effectiveness of the translation process. Detailed documentation for the usage of these commands can be found in the CodeScribe repository (Dhruv, 2024a); here, we focus on providing an overview of their functionality.
@@ 4.1. Index
This command generates a comprehensive mapping of the source tree by analyzing the project’s file hierarchy, dependencies, and code relationships. It creates a YAML file within each subdirectory, capturing critical metadata such as the relative directory location, file names, and associated constructs like modules, subroutines, and functions. This structured index is essential for navigating large legacy codebases and ensures that subsequent translation steps are both efficient and accurate. The design of the Index command embodies RAG principles by enabling efficient retrieval of relevant metadata, which aids in informed querying during later commands. Figure 4 illustrates an example of the YAML index file. The file functions as a dictionary, where the key directory stores the relative path of a subdirectory to the project root defined by the key root. The key files contains a hierarchy of dictionaries with keys corresponding to filenames and constructs. For example: [files][file1.f90][modules] provides a list of modules in the file file1.f90. During the execution of other CodeScribe commands, YAML files from all directories are loaded and compiled into an inverse dictionary that maps constructs back to their corresponding filenames for efficient querying. For instance, during draft code generation, the inverse dictionary is utilized to determine that subroutineA belongs to file2.f90.
@@ 4.4. Translate
The core functionality of CodeScribe lies in its AI-powered translation. This command transforms Fortran source code into C++ while generating the necessary Fortran-C interfaces for smooth integration between legacy and modern codebases. The translation leverages generative AI to interpret Fortran’s syntax and semantics. The seed prompt from the chat template is appended with code from the source and draft files and sent to the AI for chat completion. The source code from the Fortran file is supplied within the elements <source>...</source>, and the draft code is enclosed within <draft>...</draft>. The LLM is instructed to produce output for the C++ source and the corresponding Fortran-C interface, enclosing them within <csource>...</csource> and <fsource>...</fsource>. This mechanism is based on the implementation by the Slack team in their recent article (Slack, 2024), allowing for source code extraction and saving to separate files named <filename>.cpp and <filename>_fi.f90.
@@
5. Target Codes
We explored the application of CodeScribe to the following scenarios within scientific computing:
@@ 5. Target Codes
Developing C++ interface for Noah-MP: Noah-MP (He, C., et al., 2023) is a Fortran library that models urban canopy parameterization for atmospheric simulations. Integrating it with the Energy Research and Forecasting (ERF) Model (Almgren et al., 2023) is essential for enhancing the latter’s applicability. However, ERF is written in C++. We utilized CodeScribe to create interfaces for this integration.
LLM performance studies discussed in section 6 are dominated by MCFM code conversion because that was not only the motivator for the development of CodeScribe but has seen its most extensive use. As of this writing the entire code has been converted to C++. The conversion process initially averaged 2-3 files per day before the development of the tool. After the tool was implemented, developer productivity increased to 10-12 files per day. Below are a few additional statistics gleaned from this work.
6. LLM Performance Studies
We present two distinct studies: (1) The performance of CodeScribe with various models of different parameter counts, and (2) The influence of RAG on performance.
@@ 6.1. Model Sensitivity Study
[…]
@@
GPT-4o and GPT-3.5 Turbo, developed by OpenAI, are general-purpose models accessible via API endpoints, requiring the purchase of API credits. GPT-4o is a multimodal model capable of processing both text and images, and although OpenAI has not disclosed its architecture or parameter count, it is widely speculated to have significantly more parameters than other models. GPT-3.5 Turbo is smaller and optimized for cost-effective, high-speed text generation. While GPT-4o is well-suited for complex multimodal tasks and excels in multi-step reasoning, CodeLlama and Mistral are specialized for coding tasks and do not support multi-step reasoning to the same extent as GPT-4o. This distinction is critical when evaluating model performance in specific use cases, such as the MCFM code conversion, where the ability to engage in multi-step reasoning can significantly affect the quality of the translated code.
@@
One of the primary challenges in this translation process is the handling of array indexing. Fortran uses 1-based indexing, while C++ utilizes 0-based indexing, which can be particularly tricky when dealing with complex data structures and indirect addressing schemes such as Compressed Sparse Row (CSR) indexing. Converting between 1-based and 0-based indexing is one of the most error-prone aspects of translating between these languages, especially when it impacts index values and array bounds.
To address this challenge, we developed a set of custom container classes—FArray1D, FArray2D, FArray3D, and so on. These classes cast C++ pointers to arrays with 1-based indexing, effectively emulating the behavior of Fortran-style arrays. The container classes are designed to allow access patterns that are very similar between Fortran and C++, and they are accessed via a special operator, ensuring a consistent array handling experience across both languages. While the initialization of arrays differs between Fortran and C++, the core array management remains consistent.
For a detailed explanation of these container classes and their usage, please refer to the raw source code repository available in the supplementary materials. Specifically, the container classes are located in:
An example of their application can be found in:
(lines 56–57). These resources demonstrate the approach in full and offer insight into the indexing strategies employed to maintain consistency between Fortran and C++.
@@ 6.2. RAG Sensitivity Study
In this study, we focus on how retrieval-augmented generation (RAG) enhances code translation performance. RAG has previously been shown to improve the accuracy of natural language processing (NLP) tasks, as demonstrated by Lewis et al. (Lewis et al., 2020). To evaluate its effectiveness in code translation, we applied RAG elements within CodeScribe, specifically during the Draft command. In this step, a target.scribe file is created, and RAG is employed by parsing YAML index files and annotating the draft with prompts that provide context to the LLM, particularly about the presence of external functions.
@@ 7. Ongoing Work
In our ongoing work, we are applying the code conversion techniques developed for MCFM to design a Fortran-C interoperability layer for Noah-MP and ERF, with the goal of incorporating Urban Canopy parameterization into ERF simulations. Use of CodeScribe for Noah-MP demonstrates a systematic approach to type conversion from Fortran to C++ through modern interoperability features. The code leverages Fortran derived types defined within modular structures using constructs introduced in Fortran 2003—with selective use of Fortran 2008 features—and interfaces them with C++ via the iso_c_binding intrinsic. This guarantees that memory layouts remain consistent between the languages. In C++, mirror structures are created to directly reference Fortran-managed memory, with the implementation relying on features available from C++11 onwards. The conversion process, carried out through dedicated initialization routines, maps Fortran arrays and pointers to their C++ counterparts while preserving Fortran’s indexing conventions, thereby supporting robust and efficient cross-language data sharing. Below in an example of Noah-MP type conversion:
The example demonstrates how NoahmpIO_type is mirrored between Fortran and C++ through following implementation:
To enhance developer productivity during code conversion, we plan to expand the Inspect command to automatically generate seed_prompt.toml chat templates and to integrate CodeScribe with the LASSI framework (Dearing et al., 2024) to improve the correctness of the generated code.
@@ 8. Conclusions
In this study we presented CodeScribe, a GenAI-driven tool designed to facilitate the translation of legacy Fortran code to modern C++ while creating necessary interfaces for seamless integration. Our evaluations demonstrate that CodeScribe reduces the developer time required for code conversion and testing, showcasing its potential to enhance productivity in scientific computing projects.
@@
In addition to the MCFM code conversion, our ongoing efforts focus on applying CodeScribe to meet developer requirements across various scientific computing projects, such as Noah-MP and Flash-X, by enhancing existing commands and introducing new functionality. Flash-X (Anshu Dubey, et al., 2022) is an open-source multiphysics simulation software designed for modeling incompressible multiphase fluid dynamics with heat transfer. It is widely used in the simulation of pool and flow boiling phenomena, particularly in phase-change cooling systems (Dhruv, 2024c). Flash-X incorporates AMReX (Zhang et al., 2019) for Adaptive Mesh Refinement (AMR) grid support. However, it currently only supports the AMReX Fortran interface, which limits its compatibility with GPUs. Using CodeScribe to build GPU compatibility between Flash-X and AMReX would significantly enhance the performance and scalability of these simulations.
As we continue to refine CodeScribe and explore its applications, our aim is to further streamline the translation process and reduce overhead associated with code review and testing by exploring its integration with the LASSI framework developed at Argonne. We envision CodeScribe as a valuable tool that empowers developers in scientific computing to leverage GenAI effectively.